Submitted to the International Europhysics Conference on High Energy Physics HEP99, Tampere, Finland, July 1999

your wonderful title ... at HERA

H1 Collaboration

Abstract

...your striking abstract

Abstract: Parallel sessions: Plenary sessions:

Electronic Access: http://www-h1.desy.de/publications/

1 Introduction

...your clearly written text with figures (such as Fig.1, from [1]) and important relations

x_{IP} = \frac{Q^{2} + M_{X}^{2} − t}{Q^{2} + W^{2} + M_{p}^{2}} .  (1)

Equation (1) gives ...

Figure 1: The average squared transverse momentum ⟨p_{t}^{∗2}⟩ of hadrons is shown as a function of Feynman x_{F}.

Acknowledgements References

[1] H1 Collab., C. Adloff et al., Phys. Lett. B 428 (1998) 206